OpenStudy (anonymous):

Brooks built two similar wooden chests in the shape of rectangular prisms. She used 5 pints of paint to cover the outside of the larger chest fully and 2 pints to cover the outside of the smaller chest. How much more volume does the larger chest contain? A. 158% more B. 250% more C. 300% more D. 395% more

OpenStudy (anonymous):

Even though I have already solved this problem, I will post how I received my answer in case anyone ever needs help with the same or a similar one to this... 1. Find the ratio of the similar solids: ratio of chests is 5/2 2. Find the square root of the ratio: √5/2 = 1.58113883008 (or 1.58 when rounded) 3. Now, cube your answer: 1.58113883008(cubed) = 3.95284707521 (or 3.95 when rounded) 4. Convert decimal into a percentage: 3.95 = 395%